A circular flower bed is 7 meters in radius. Santan bushes are to be planted around the edge. If they are to be about half a meter apart, how many bushes are to be planted?: A circular flower bed is 7 meters in radius. Santan bushes are to be planted around the edge. If they are to be about half a meter apart, how many bushes are to be planted?
Answer by checkley77(3654)

C=2*PI*R
C=2*3.14*7
C=43.96 IS THE CIRCUMFERANCE.
43.96/2=21.98 PLANTS OR ROUNDED 22 PLANTS WILL DO.